Hello!
Mel plans to fence his yard for his new puppy. The yard is a 41 ft by 125 ft rectangle. Fencing costs $15 per 10 ft section. What is the cost of the fence?
I always thought I was good at word problems but I cant seem to get this one started. Even if i can get help with just starting this off I'm sure I'll figure it out. Thank you so much for the help!

Answer by ewatrrr(24785)


Hi

P = 2*125ft + 2*41ft = 332ft (332/10)*$15 = $498